# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 7. Weight function
$\frac{16\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The integration over $D_{1}$ has the following upper bound, when both of the terms of Expression (7.5) are added together:
$\frac{1}{\sqrt{\sinh(\rho)}}\left(12+16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\right)$
$\leq\ 16\pi\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
Next, we will focus on bounding the integral over $D_{2}$. We write it using the polar coordinates around $z$. Thus $x=\pi_{1}(\phi_{R}(z,\theta))$, where $\theta=0$ corresponds to direction $z,z^{\prime}](z)$. Then the integral equals:
$\int_{-\Psi}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta$
= $=\ 2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta,$
where the equality followed from using the symmetry. $r=r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates $R$ and $\theta$. $R(\theta)$ denotes the minimum distance that needs to be propagated into direction $\theta$ so that we are in $D_{2}$. Notice also that $\Psi$ is less than $\pi/2$. By using the hyperbolic law of cosines we have that
$\cosh(t^{\prime})=\cosh(\rho)\cosh(t)-\sinh(\rho)\sinh(t)\cos(\Psi),$
$\cosh(r)=\cosh(\rho)\cosh(R)-\sinh(\rho)\sinh(R)\cos(\theta).$
As $\cosh(r)$ grows as $R$ grows, we have the following lower bound for $\cosh(t^{\prime})-\cosh(r)$:
$\sinh(\rho)\sinh(t)(\cos(\theta)-\cos(\Psi)).$
@@
Hence the integral expression has the following upper bound:
$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\frac{1}{\sqrt{\sinh(t)}}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}}dRd\theta.$
This has an upper bound when we integrate from 0 instead of $R(\theta)$; the following expression is acquired after we evaluate the integral:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}d\theta.$
This is then bounded by
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}d\theta,$
which has the following upper bound using the mean value theorem:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
Using the well-known bound for sine, we get the following upper bound:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\frac{2}{\pi}\theta}}d\theta.$
This is actually the Beta integral (see [Olv10, Chp. 5]), so we get the following equal expression:
$\frac{4\sqrt{\pi}\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}}.$
The integration over both $D_{1}$ and $D_{2}$ is done. Uniting the bounds of these two cases gives us the final bound for $F_{t,t^{\prime},\rho}$:
@@
$\frac{1}{\sqrt{\sinh(\rho)}}\left(16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\ +\ \frac{4\pi\sqrt{\pi}}{\sqrt{2}}\right)$
@@
$\leq\qquad$ $\frac{32\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The following lemma takes care of the last case, see Figure 4.
@@
where
$C_{4}=99.$
The proof is again based on splitting the integration domain into two. This time the splitting is done by $w,w^{\prime}]$ which splits the integration into two sides, one closer to $z$ and one closer to $z^{\prime}$. The technique used to bound both sides is the same, based on establising carefully bounds using hyperbolic geometry.
@@
The integration domain of $F_{t,t^{\prime},\rho}$ is the intersection of $B(z,t)$ and $B(z^{\prime},t^{\prime})$. Since $t<\rho+t^{\prime}$, the intersection is not a ball. Let $w$ and $w^{\prime}$ be the intersections of the boundaries of the balls, and let $m$ be the intersection of $z,z^{\prime}]$ with $w,w^{\prime}]$. We parametrize the domain using the polar coordinates around $m$ with $\theta=0$ corresponding to direction $z,z^{\prime}](m)$. Hence $x=\pi_{1}(\phi_{r}(m,\theta))$. To easen the notation, we call $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$. We will present $d(z,x)$ and $d(z^{\prime},x)$ using polar coordinates, denoted by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectfully. For illustration, see Figure 4. Hence $F_{t,t^{\prime},\rho}$ is in polar coordinates
@@
where $r(\theta)$ is the maximum distance we can go from $m$ to direction $\theta$ to still stay in the integration domain. We can split the integration into two parts:
@@
The first term integrates over the side closer to $z^{\prime}$ while the second term integrates the side closer to $z$, these sides are separated by $w,w^{\prime}]$.
Let us focus on the integrand of the first term. Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where we note that $\cos(\theta)$ is positive. We also notice that by using the hyperbolic law of cosines we get that
@@
Inserting the earlier equality of $\cosh(t)$ into the previous yields the following:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\sinh(a)\left(\sinh(r(\theta))-\sinh(r)\right),$
which then has a lower bound:
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
and then has a lower bound
$\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))$ is bounded from below by
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
This equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(b)\sinh(a)\sinh(r(\theta))+\cosh(a)\sinh(b)\sinh(r)\right),$
and is then bounded from below by
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(b)\sinh(a)+\cosh(a)\sinh(b)\right).$
As $a+b=\rho$, the previous equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
Now, let us focus on the integrand of the second term of Expression (7.14). Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))-\sinh(b)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where $\cos(\theta)$ is negative. We also notice that, using the hyperbolic law of cosines,
$\cosh(t)=\cosh(a)\cosh(h),$
$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
which implies that
$\cosh(t)=\frac{\cosh(a)}{\cosh(b)}\cosh(t^{\prime}).$
The previous gives the following equality after inserting an earlier equality of $\cosh(t^{\prime})$ in:
$\cosh(t)=\cosh(a)\cosh(r(\theta))-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)-\cos(\theta)\sinh(b)\left(\sinh(r(\theta))+\sinh(r)\right).$
This has a lower bound
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))-\sinh(a)\sinh(r)\right),$
and then has a lower bound
$(-1)\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R(r,\theta)))(\cosh(t^{\prime})-\cosh(P(r,\theta)))$ is bounded from below by
$(-1)\,\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)\,\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
This equals
$(-1)\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(a)\sinh(b)\sinh(r(\theta))+\cosh(b)\sinh(a)\sinh(r)\right),$
and is then bounded from below by
$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(a)\sinh(b)+\cosh(b)\sinh(a)\right).$
As $a+b=\rho$, the previous equals
@@
These observations for the integrands of the terms of Expression (7.14) imply that we get the following upper bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta$ (7.15)
Focus on the inner integral. Assume first that $r(\theta)<2$. The integral is bounded by the following by using the mean value theorem:
$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}.$
This is bounded by $2\sqrt{2}$. Then, assume that $r(\theta)>2$. We split the integration as follows:
$\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$
$a$ is specified later. We have the following upper bound for the previous:
$\frac{1}{\sqrt{\sinh(a)}}\,\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr.$
The first term was bounded using the boundary of the integration domain and the second term was bounded using the mean value theorem. By evaluating the integrals, we acquire the following equal expression:
$\frac{2\,\sqrt{\cosh(r(\theta))-\cosh(a)}}{\sqrt{\sinh(a)}}\ +\ 2\,\sqrt{r(\theta)-a}\leq\frac{2\,\sqrt{r(\theta)-a}\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta))}}\ +\ 2\,\sqrt{r(\theta)-a}.$
The inequality followed from using the mean value theorem. We choose $a$ to equal $r(\theta)-1$. Then the previous equals
$\frac{2\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta)-1)}}\ +\ 2\leq\frac{2\,\sqrt{\sinh(2)}}{\sqrt{\sinh(1)}}\ +\ 2,$
where for the inequality we used that $r(\theta)>2$. The previous is bounded by $9$.
The previous observations mean that Expression (7.15) has the following upper bound:
$\frac{9}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta.$
The integral is bounded from above by $11$, so we can conclude that
$F_{t,t\,\rho}\leq\frac{99}{\sqrt{\sinh(\rho)}}.$

### v2

@@ 7. Weight function
$2\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\left(\int_{\Theta}^{\pi}\frac{d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}}\right)dr,$ (7.7)
where only half of the angles are considered due to symmetry, giving the coefficient $2$ in front. Using the hyperbolic law of cosines, we record two useful equalities:
$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})+\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.8)
$\cosh(R(r,\theta))=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$ (7.9)
Focus on the inner integral of Expression (7.7). First, we write it using Expressions (7.8) and (7.9):
$\int_{\Theta}^{\pi}\frac{d\theta}{\sqrt{\cosh(\rho)\cosh(t^{\prime})+\sinh(\rho)\sinh(t^{\prime})\cos(\Theta)-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta)}}.$
We then split it into three terms by dividing the integration domain:
$\int_{\Theta}^{\Theta+\theta_{0}}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}$
$+\ \int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}$
$+\ \int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}.$ (7.10)
We remark that $\Theta\in(0,\frac{\pi}{2})$ since $z^{\prime}$ is contained in $B(z,t)$ as $\rho\in(t-t^{\prime},t)$. $\theta_{0}\in(0,\frac{\pi}{2}-\Theta)$ and will be specified later. We note that $\theta_{0}$ can depend on $r$. We bound these terms separately.
We start by bounding the first term of Expression (7.10). It is bounded from above by
$\int_{\Theta}^{\Theta+\theta_{0}}\frac{d\theta}{\sqrt{\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))}}=\frac{\theta_{0}}{\sqrt{\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))}},$
since $\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))$ is positive.
Next, we bound the second term of Expression (7.10). It is bounded from above by
$\int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}\sqrt{\cos(\Theta)-\cos(\theta)}},$
since $\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))$ is positive and $t^{\prime}>r$. By examining the behavior of the function, we notice that
$\cos(\Theta)-\cos(\theta)\geq 1-\cos(\theta-\Theta),\qquad\theta\in(\Theta,\tfrac{\pi}{2}).$
@@
Using this gives the following upper bound for the previous integral:
$\frac{1}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}}\int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-\cos(\theta-\Theta)}}\leq\frac{1}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}}\frac{\frac{\pi}{2}}{\sqrt{1-\cos(\theta_{0})}}.$
This is bounded from above by
$\frac{\pi^{2}}{2\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\frac{1}{\theta_{0}}$
using
$1-\cos(x)\geq\frac{2}{\pi^{2}}x^{2},\qquad x\in(-\pi,\pi).$
Finally, we bound the third term of Expression (7.10). It is bounded from above by
$\int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})\,(-\cos(\theta))}}\leq\frac{3}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}},$
since $\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))$ and $\sinh(\rho)\sinh(t^{\prime})\cos(\Theta)$ are positive. The inequality is obtained by integrating $\theta$ and then bounding the result from above by $3$.
The three terms of Expression (7.10) are bounded. Hence we have an upper bound for Expression (7.10), and thus also for Expression (7.7), which is:
$\frac{2}{\sqrt{\cosh(\rho)}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,\theta_{0}\,dr}{\cosh(t^{\prime})-\cosh(r)}+\frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\theta_{0}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
$+\ \frac{6}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
We choose
@@
$\theta_{0}=\min\left(\frac{\pi}{2}-\Theta,\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\right),$
with $\alpha\in(0,\frac{1}{2})$. The previous integral is then bounded from above by
$\frac{2}{\sqrt{\cosh(\rho)}\sinh(t^{\prime})^{\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{1-\alpha}}$
$+\ \frac{\pi^{2}}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{\frac{1}{2}+\alpha}}$
$+\ \frac{6}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
The upper bound was obtained by replacing $\theta_{0}$ with
$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha};$
this is possible since the middle term would vanish if
@@
$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\geq\frac{\pi}{2}-\Theta.$
Evaluating the integrals gives
$\frac{2}{\alpha}\,\frac{(\cosh(t^{\prime})-1)^{\alpha}}{\sinh(t^{\prime})^{\alpha}}\,\frac{1}{\sqrt{\cosh(\rho)}}+\frac{\pi^{2}}{\sqrt{2}(\frac{1}{2}-\alpha)}\,\frac{(\cosh(t^{\prime})-1)^{\frac{1}{2}-\alpha}}{\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\,\frac{1}{\sqrt{\sinh(\rho)}}+6\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\,\frac{1}{\sqrt{\sinh(\rho)}}.$
Choosing $\alpha=\frac{1}{4}$ yields the upper bound:
$\frac{42}{\sqrt{\sinh(\rho)}}.$
This finishes the bound for the integration over $D_{1}$ in Expression (7.6).
We proceed to bound the integration over $D_{2}$ in Expression (7.6). We write $x$ using polar coordinates around $z$, taking $\theta=0$ in the direction $z,z^{\prime}](z)$. Hence $x=\pi_{1}(\varphi_{R}(z,\theta))$. $r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates. We notice that the integration domain of the angles is $(-\Psi,\Psi)$, where $\Psi$ is the angle between $z,z^{\prime}](z)$ and $[z,w](z)$. Furthermore, the integration domain of the distances depends on the angle: $R\in(R(\theta),t)$, where $R(\theta)$ is the minimum distance required to reach the integration domain in the direction $\theta$. Figure 5 illustrates this.
Figure 5. The integration domain $D_{2}$ consists of points in the area surrounded by the segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ and the arc between $w$ and $w^{\prime}$ on $\partial B(z,t)$.
Hence the second term of Expression (7.6) equals
$2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r(R,\theta))}}dR\,d\theta,$ (7.11)
where only half of the angles are considered due to symmetry, giving the coefficient $2$ in front. Using the hyperbolic law of cosines, we record two useful equalities:
$\cosh(t^{\prime})=\cosh(t)\cosh(\rho)-\sinh(t)\sinh(\rho)\cos(\Psi),$
$\cosh(r(R,\theta))=\cosh(R)\cosh(\rho)-\sinh(R)\sinh(\rho)\cos(\theta).$
We remark that $r(R,\theta)$ grows as $R$ grows. Hence we have the following inequality using the previous equalities:
$\cosh(t^{\prime})-\cosh(r(R,\theta))\geq\sinh(t)\,\sinh(\rho)\,(\cos(\theta)-\cos(\Psi)).$ (7.12)
Focus on the inner integral of Expression (7.11). We acquire an upper bound by using the inequality of Expression (7.12):
$\int_{R(\theta)}^{t}\frac{\sinh(R)\,dR}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\sinh(\rho)\,\sinh(t)\,(\cos(\theta)-\cos(\Psi))}}$
$\leq\ \frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}\,\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{1}{\sqrt{\sinh(\rho)}}$
$\leq\ \frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{1}{\sqrt{\sinh(\rho)}},$
where the first inequality is obtained by evaluating the integral over the enlarged domain $(0,t)$.
We now have a bound for the inner integral of Expression (7.11), giving the following upper bound for the whole expression:
$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\cos(\theta)-\cos(\Psi)}}\leq\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
The inequality follows from the mean value theorem. The previous expression is further bounded by
$\sqrt{2}\sqrt{\pi}\,\frac{1}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\theta}\,\sqrt{\Psi-\theta}},$
using that
$\sin(x)\geq\frac{2}{\pi}x,\qquad x\in(-\frac{\pi}{2},\frac{\pi}{2}).$
The integral is Euler’s beta integral and equals $\pi$, see [LOZ03, Chp. 5.12.]. Hence the expression equals
$\sqrt{2}\pi\sqrt{\pi}\,\frac{1}{\sqrt{\sinh(\rho)}}.$
We have now bounded the integration over $D_{2}$ in Expression (7.6).
Combining the bounds for $D_{1}$ and $D_{2}$, we get the final upper bound for $F_{t,t^{\prime},\rho}$:
$(42+\sqrt{2}\pi\sqrt{\pi})\,\frac{1}{\sqrt{\sinh(\rho)}}\leq 50\,\frac{1}{\sqrt{\sinh(\rho)}}.$
This concludes the proof of Lemma 7.5.
We state the third lemma concerning the case of $\rho\in(t,t+t^{\prime})$.
@@
where $C_{4}=38$.
The following figure, Figure 6, illustrates the integration domain.
Figure 6. The case of $\rho\in(t,t+t^{\prime})$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball. The integration domain is split into two parts by $w,w^{\prime}]$. Notice that $a+b=\rho$, the separation of $z$ and $z^{\prime}$.
The proof is again based on splitting the integration domain into two. This time the splitting is done by $w,w^{\prime}]$, which separates the domain into two sides: one closer to $z$ and one closer to $z^{\prime}$. Here, $w$ and $w^{\prime}$ are points of $\partial B(z,t)\cap\partial B(z^{\prime},t^{\prime})$. The technique used to bound both sides is the same and relies on hyperbolic geometry, in particular the hyperbolic law of cosines.
@@
The integration domain of $F_{t,t^{\prime},\rho}$ is the intersection of $B(z,t)$ and $B(z^{\prime},t^{\prime})$. Since $t<\rho+t^{\prime}$, the intersection is not a ball. Let $w$ and $w^{\prime}$ be the intersections of the boundaries of the balls, and let $m$ be the intersection of $z,z^{\prime}]$ with $w,w^{\prime}]$. We parametrize the domain using polar coordinates around $m$, with $\theta=0$ corresponding to the direction $z,z^{\prime}](m)$. Hence $x=\pi_{1}(\phi_{r}(m,\theta))$. To ease notation, we let $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$. We denote $d(z,x)$ and $d(z^{\prime},x)$ in polar coordinates by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectively. Then $F_{t,t^{\prime},\rho}$ can be written as
@@
where $r(\theta)$ is the maximal distance from $m$ in direction $\theta$ within the integration domain. We split the integration into two parts:
@@
The first term integrates over the side closer to $z^{\prime}$, while the second integrates over the side closer to $z$; these sides are separated by $w,w^{\prime}]$.
For the first term of Expression (7.14), the hyperbolic law of cosines gives:
$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$ (7.15)
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$ (7.16)
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$ (7.17)
where $\cos(\theta)>0$. Also, we have
@@
Inserting Equality (7.15) gives
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$ (7.18)
Lower bounds for the differences are
$\cosh(t)-\cosh(R)\geq\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right),$
$\cosh(t^{\prime})-\cosh(P)\geq\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
and their product satisfies
$(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))\geq\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
For the second term of Expression (7.14), using the same approach (the roles of $t,R$ and $t^{\prime},P$ change), we have the following lower bound:
@@
Notice that 1-1 comes from the fact that $\cos(x)$ is negative when $x\in(\tfrac{\pi}{2},\tfrac{3\pi}{2})$. These bounds give the following upper bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta.$ (7.19)
For the inner integral, if $r(\theta)<2$, the mean value theorem gives the following upper bound:
$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}\leq 2\sqrt{2}.$
If $r(\theta)>2$, we split the inner integral of Expression (7.19) at $c\in(0,r(\theta))$ to write it as:
$\int_{0}^{c}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{c}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$ (7.20)
c is specified later.
The first term of the previous expression has the following upper bound using the boundary of the integration domain:
$\frac{1}{\sqrt{\cosh(r(\theta))-\cosh(c)}}\int_{0}^{c}\frac{e^{\tfrac{r}{2}}}{\sqrt{2}}\,dr$ $=\frac{2}{\sqrt{2}}\frac{e^{\tfrac{c}{2}}-1}{\sqrt{\cosh(r(\theta))-\cosh(c)}}\leq\sqrt{2}\frac{\sinh\!\left(\tfrac{c}{2}\right)}{\sqrt{\sinh(c)\,(r(\theta)-c)}},$
where the inequality follows from the mean value theorem together with the bound for $\sinh$. Since
$\frac{\sinh\!\left(\tfrac{x}{2}\right)}{\sqrt{\sinh(x)}}\leq 1,$
we conclude that the first term of Expression (7.20) is bounded from above by
$\frac{\sqrt{2}}{\sqrt{r(\theta)-c}}.$
The second term of Expression (7.20) is bounded from above by the following expression using the mean value theorem:
$\int_{c}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)-c}.$
With both terms bounded, Expression (7.20) has the following upper bound by choosing $c=r(\theta)-1$:
$\frac{\sqrt{2}}{\sqrt{r(\theta)-c}}+2\sqrt{r(\theta)-c}=2+\sqrt{2}.$
Hence, Expression (7.19) is bounded by
$\frac{2+\sqrt{2}}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta\leq\frac{38}{\sqrt{\sinh(\rho)}},$
where the last integral is bounded from above by $11$. This completes the proof of Lemma 7.6.